OpenStudy (anonymous):

Multiply. Write your answer in lowest terms. 9/22(- 2/9)

OpenStudy (btaylor):

When you multiply fractions, multiply the numerators and the denominators: \[\frac{9}{22} \times \frac{-2}{9} \rightarrow \frac{9 \times -2}{22 \times 9}\] Before you multiply it out fully though, see what cancels out. Here, there is a 9 on top and bottom, so that can be eliminated. Also, 2 and 22 are both divisible by 2, so simplify that: \[\frac{9 \times -2}{22 \times 9} \rightarrow \frac{\cancel{9} \times -2}{22 \times \cancel{9}} \rightarrow \frac{-2}{22} \rightarrow \frac{\cancel{2} \times -1}{\cancel{2} \times 11} \rightarrow -\frac{1}{11}\]